# Diffusion-Driven Universal Model Inversion Attack for Face Recognition

### v1

@@ V-C Out-Of-Domain Detection
TABLE V: Out-of-domain detection rate (%) across models and datasets.
@@
V-D Limited Impact of Prior Knowledge on Generator Training
We assess whether incorporating prior knowledge into the generator enhances reconstruction performance. To this end, we train a DDPM on CelebA-HQ and attack target identities also drawn from CelebA-HQ (with disjoint identities), ensuring that the generator and target images share the same distribution. Results are compared against attacks on the LFW dataset. As shown in Tab. III and Tab. V, it is surprising that performance varies inconsistently between CelebA-HQ and LFW across both privacy attacks and OODD tasks. These findings indicate that, in training-free, open-set model inversion, prior knowledge of the data distribution has minimal impact on reconstruction accuracy. Even with shared distributions, the generator cannot reliably synthesize identity-consistent images without a well-defined identity-guided mechanism.
VI Ablation Study
We conduct ablation studies on the proposed DiffUMI and OODD using the LFW [35] and ImageNet [38] datasets, with the privacy-preserving PartialFace [6] model as the target. Type II accuracy is averaged across all four face recognition models for joint evaluation, as outlined in Sec. IV. In each study, we vary a single parameter while holding all others constant to isolate its impact on performance, with parameter settings provided in Tab. II.
VI-A Model Inversion (DiffUMI)
Reliability of Latent Codes. We propose a latent code generation strategy in Sec. II-D that combines D’Agostino’s $K^{2}$ test [40, 41, 42] with MTCNN [43] to improve latent code reliability. As shown in Tab. VI, our approach achieves the highest latent code reliability, leading to superior performance.
Top N Selection. As detailed in Sec. II-E, DiffUMI does not necessitate processing all latent codes. While optimal initialization typically improves reconstruction accuracy, it does not always yield the best results after manipulation. Consequently, we adopt a top $N$ selection strategy, where increasing $N$ enhances reconstruction accuracy at the expense of higher computational cost. Based on our evaluation, we recommend $N=3$ as an optimal trade-off, achieving strong attack performance with manageable computational overhead, as shown in Tab. VII.
TABLE VI: Performance across different latent code reliabilities.
@@ VI-A Model Inversion (DiffUMI)
TABLE VII: Performance (%) for varying values of $N$.
@@
Ranked Adversary. We propose the Ranked Adversary strategy in Sec. II-F to efficiently identify a satisfactory, rather than optimal, solution for model inversion. By prioritizing sequential initialization and allowing early termination once a predefined condition is met, this strategy reduces unnecessary computations. As shown in Tab. VIII, it significantly lowers computational cost with only a minor trade-off in Type II accuracy and OODD performance, effectively balancing efficiency and attack performance.
TABLE VIII: Performance (%) with and without the ranked adversary strategy. “NO” indicates that all selected $N$ latent codes are processed, and the reconstruction with the highest similarity is selected as the final output.
@@
Attack Threshold. We define the attack threshold $\tau_{A}$ in Sec. II-F1 as the criterion for successful manipulation, where the reconstructed image achieves sufficient similarity to the target. This threshold also serves as the early termination criterion in the ranked adversary strategy. $\tau_{A}$ must satisfy $\tau_{A}\gg\tau_{F}$ for robustness. However, excessively high values of $\tau_{A}$ (e.g., $\tau_{A}=1$) lead to unnecessary computational cost and the risk of overfitting.
For the target model, PartialFace [6], used in this ablation study, the maximum embedding similarity achievable by real facial images is 0.98. As shown in Tab. IX, $\tau_{A}<0.98$ results in faster model inversion completion, whereas $\tau_{A}>0.98$ improves Type II accuracy at the cost of increased computational overhead. Additionally, higher $\tau_{A}$ values increase the risk of being detected as out-of-domain due to greater distortion in the reconstructed images. These results demonstrate that our proposed strategy (Sec. II-F1), which defines $\tau_{A}$ as the maximum embedding similarity achievable by real facial images in the target model, is optimal.
TABLE IX: Performance (%) across varying attack threshold $\tau_{A}$.
@@
Perturbation Constraint. As discussed in Fig. 4, the $L_{2}$-norm constrained adversary better preserves normality compared to the $L_{\infty}$-norm adversary, leading to improved reconstruction fidelity. This is further validated by Tab. X, where $L_{2}$-norm ($\epsilon=35$) and $L_{\infty}$-norm ($\epsilon=0.15$) achieve similar similarity values, yet the $L_{2}$-norm attack demonstrates significantly superior performance. Additionally, Tab. X and Fig. 12 show that while increasing the attack magnitude $\epsilon$ enhances Type II accuracy and accelerates the attack, it also introduces more artifacts, raising the likelihood of detection as out-of-domain inputs.
TABLE X: Performance (%) under varying perturbation constraints.
@@
VI-B Out-of-Domain Detection (OODD)
We propose OODD framework in Sec. III, integrating two failure cases: matching failure and face detection failure. Both criteria must be considered together, as relying on a single strategy alone is insufficient for effective OODD. As shown in Tab. XI (see red values), using only one strategy results in an OODD rate of approximately 50% for out-of-domain inputs (i.e., ImageNet), highlighting its limitations.
TABLE XI: OODD rate (%) when jointly considering both failure cases (ours) versus individual strategies.
@@ VI-B Out-of-Domain Detection (OODD)
VII Discussion on Defenses Against DiffUMI
We have demonstrated the success of our attack on two methods designed to mitigate model inversion by reducing facial information exposure, either through randomization of frequency components or by managing privacy budgets [5, 6]. However, more advanced strategies, such as federated learning, aim to counter privacy recovery. Investigating whether our method can generalize to these broader defense strategies represents a valuable extension of our work.
Nevertheless, the core strategy of our model inversion attack is to refine facial reconstruction while minimizing artifacts to avoid overfitting to the target model. We argue that if a facial image can match the target embedding without overfitting due to artifacts, it must correspond to the correct identity. This optimization approach relies solely on the input and output of the target model, which cannot bypass in an application context. Therefore, we believe our model inversion attack should be effective against any target system providing input-output access, regardless of integrated defense mechanisms.
We posit that the most efficient defense against our attack would involve hiding gradients to prevent white-box model inversion. However, while our black-box version is computationally more intensive, it is not impossible to execute, though it requires significantly more time. We acknowledge this as a limitation and note that further testing is needed to verify these discussions in future work.
VIII Conclusion, Limitation, and Future Work
@@
Despite these contributions, we acknowledge three limitations. First, the computational cost of black-box DiffUMI remains prohibitively high, limiting its practical deployment. Second, modifying latent codes without introducing significant artifacts using diffusion-driven unconditional image generation presents a fundamental challenge, restricting the extent of feasible modifications. Third, further investigation is needed to uncover the mathematical reasons behind the poor performance of coarse-grained adversarial manipulation in unconditional image generation.
@@
For future work, beyond addressing the aforementioned limitations, we aim to extend DiffUMI to broader domains, including classification models for datasets such as CIFAR-10, ImageNet, and medical imaging. Additionally, we intend to evaluate the potential defenses against our model inversion attack, as outlined in Sec. VII. We also plan to develop objective evaluation metrics to replace human assessment, as subjective judgments, such as “Do these two images belong to the same person?”, are inherently inconsistent.
In extending OODD, we would like to investigate its potential applications in data poisoning detection and defense against malicious inputs. Additionally, we consider developing a novel OODD framework independent of model inversion to enhance its applicability and robustness.

### v2

@@ 5.3 Out-Of-Domain Detection
TABLE VI: Out-of-domain detection rate (%) across models and datasets.
@@
5.4 Prior Knowledge on Generator Training
We assess whether incorporating prior knowledge into the generator enhances reconstruction performance. To this end, we train a DDPM on CelebA-HQ and attack target identities also drawn from CelebA-HQ (with disjoint identities), ensuring that the generator and target images share the same distribution. Results are compared against attacks on the LFW dataset. As shown in Tab. III, Tab. IV and Tab. VI, it is surprising that performance varies inconsistently between CelebA-HQ and LFW across both privacy attacks and OODD tasks. These findings indicate that, in training-free, open-set model inversion, prior knowledge of the data distribution has minimal impact on reconstruction accuracy. Even with shared distributions, the generator cannot reliably synthesize identity-consistent images without a well-defined identity-guided mechanism.
6 Ablation Study
We conduct ablation studies on the proposed DiffUMI and OODD using the LFW [40] and ImageNet [41] datasets, with the privacy-preserving PartialFace [7] model as the target. Type II accuracy is averaged across all four face recognition models for joint evaluation, as outlined in Sec. 4. In each study, we vary a single parameter while holding all others constant to isolate its impact on performance, with parameter settings provided in Tab. II.
6.1 Model Inversion (DiffUMI)
Reliability of Latent Codes. We propose a latent code generation strategy in Sec. 2.4, combining D’Agostino’s $K^{2}$ test with MTCNN to improve latent code reliability [43, 44, 45, 46]. As shown in Tab. VII, our approach achieves the highest latent code reliability, leading to superior performance.
Top N Selection. As detailed in Sec. 2.5, DiffUMI does not necessitate processing all latent codes. While optimal initialization typically improves reconstruction accuracy, it does not always yield the best results after manipulation. Consequently, we adopt a top $N$ selection strategy, where increasing $N$ enhances reconstruction accuracy at the expense of higher computational cost. Based on our evaluation, we recommend $N=3$ as an optimal trade-off, achieving strong attack performance with manageable computational overhead, as shown in Tab. VIII.
TABLE VII: Performance across different latent code reliabilities.
@@ 6.1 Model Inversion (DiffUMI)
TABLE VIII: Performance (%) for varying values of $N$.
@@
Ranked Adversary. We propose the Ranked Adversary strategy in Sec. 2.6 to efficiently identify a satisfactory, rather than optimal, solution for model inversion. By prioritizing sequential initialization and allowing early termination once a predefined condition is met, this strategy reduces unnecessary computations. As shown in Tab. IX, it significantly lowers computational cost with only a minor trade-off in Type II accuracy and OODD performance, effectively balancing efficiency and attack performance.
TABLE IX: Performance (%) with and without the ranked adversary strategy. “NO” indicates that all selected $N$ latent codes are processed, and the reconstruction with the highest similarity is selected as the final output.
@@
Attack Threshold. We define the attack threshold $\tau_{A}$ in Sec. 2.6.1 as the criterion for successful manipulation, where the reconstructed image achieves sufficient similarity to the target. This threshold also serves as the early termination criterion in the ranked adversary strategy. $\tau_{A}$ must satisfy $\tau_{A}\gg\tau_{F}$ for robustness. However, excessively high values of $\tau_{A}$ (e.g., $\tau_{A}=1$) lead to unnecessary computational cost and the risk of overfitting.
For the target model, PartialFace [7], used in this ablation study, the maximum embedding similarity achievable by real facial images is 0.98. As shown in Tab. X, $\tau_{A}<0.98$ results in faster model inversion, whereas $\tau_{A}>0.98$ improves Type II accuracy at the cost of increased computational overhead. Additionally, higher $\tau_{A}$ values raise the risk of being detected as out-of-domain due to greater distortion in the reconstructed images. These results demonstrate that our proposed strategy (Sec. 2.6.1), which defines $\tau_{A}$ as the maximum embedding similarity achievable by real facial images in the target model, is optimal.
TABLE X: Performance (%) across varying attack threshold $\tau_{A}$.
@@
Perturbation Constraint. As discussed in Fig. 4, the $L_{2}$-norm constrained adversary better preserves normality compared to the $L_{\infty}$-norm adversary, leading to improved reconstruction fidelity. This is further validated by Tab. XI, where $L_{2}$-norm ($\epsilon=35$) and $L_{\infty}$-norm ($\epsilon=0.15$) achieve similar similarity values, yet the $L_{2}$-norm attack demonstrates significantly superior performance. Additionally, Tab. XI and Fig. 13 show that while increasing the attack magnitude $\epsilon$ enhances Type II accuracy and accelerates the attack, it also introduces more artifacts, raising the likelihood of detection as out-of-domain inputs.
TABLE XI: Performance (%) under varying perturbation constraints.
@@
6.2 Out-of-Domain Detection (OODD)
We propose OODD framework in Sec. 3, integrating two failure cases: matching failure and face detection failure. Both criteria must be considered together, as relying on a single strategy alone is insufficient for effective OODD. As shown in Tab. XII (see red values), using only one strategy results in an OODD rate of approximately 50% for out-of-domain inputs (i.e., ImageNet), highlighting its limitations.
TABLE XII: OODD rate (%) when jointly considering both failure cases (ours) versus individual strategies.
@@ 6.2 Out-of-Domain Detection (OODD)
7 Conclusion
@@
For future work, we aim to extend DiffUMI to broader domains, including classification models for datasets such as CIFAR-10, ImageNet, and medical imaging. Additionally, we intend to evaluate the potential defenses against our model inversion attack. We also plan to develop objective evaluation metrics to replace human assessment, as subjective judgments, such as “Do these two images belong to the same person?”, are inherently inconsistent. In extending OODD, we would like to investigate its potential applications in data poisoning detection and defense against malicious inputs. Additionally, we consider developing a novel OODD framework independent of model inversion to enhance its applicability and robustness.
Appendix A Training-Dependent Attacks and Naive APGD
As shown in Tab. I, MAP2V [9] is the most recent and relevant training-free baseline for open-set face recognition. We also compare with training-dependent attacks [27, 17] in Tab. XIII and Fig. 14, though they are not primary baselines due to a key limitation: each target model requires a separately trained generator, leading to high computational cost. We exclude closed-set face recognition attacks, as they rely on class labels rather than target embeddings and thus follow different assumptions.
As shown in Tab. XIII, we evaluate two training-dependent attacks [27, 17], each using a generator trained specifically for the target model ArcFace [5]. The training (FFHQ [49]) and testing (LFW [40]) datasets are disjoint, ensuring an open-set setting. In contrast, both MAP2V [9] and our DiffUMI are training-free. Despite the advantage of training-dependent methods, DiffUMI performs comparably to DSCasConv [17] and significantly outperforms the others.
Figure 14: Visual fidelity comparison across attack methods shows that both Shahreza et al. [27] and our method produce high-resolution, headshot-style reconstructions with superior visual quality. In contrast, DSCasConv [17] yields the lowest fidelity, with reduced resolution, blurred features, and limited facial detail. Naively applying APGD to the DDPM latent space introduces severe visual artifacts, even when starting from the top-1 latent code out of 1,000 candidates.
TABLE XIII: Comparison of DiffUMI with Benchmark Attack Methods.
Attack Generator Training- Free Dataset Target Model Test Model
FaceNet ArcFace DCTDP PartialFace Avg.
Type I Type II Type I Type II Type I Type II Type I Type II Type I Type II
[…]
  • denotes direct application of APGD to the DDPM latent space without additional strategies. “Top 1/1000” refers to initialization from the best latent code among 1,000 random samples, while “Random” indicates no pre-selection (i.e., Top 1/1). Gray cells indicate settings where the target and test models are the same. Our training-free attack matches the performance of the training-dependent DSCasConv, while outperforming other baselines.
Despite DSCasConv’s good reconstruction accuracy, it faces substantial limitations. Specifically, it took 1.5 days of training on dual A100 80GB GPUs and is constrained to generating $112\times 112$ resolution images for a specific target model. Moreover, DSCasConv produces the lowest visual fidelity (see Fig. 14), with blurred features and limited facial detail. These limitations highlight the value of training-free approaches and underscore the advantages of leveraging generative models such as GANs or diffusion models.
We also compare our method with a baseline that directly applies APGD to the latent space of diffusion models. As shown in Tab. XIII and Fig. 14, this naive strategy achieves near-perfect accuracy (nearly 100%) on the target model but suffers from severe adversarial artifacts and overfitting, resulting in 11.97%–29.7% lower Type II accuracy than our method on test models. These results underscore the novelty of our approach, which manipulates unconditional diffusion generation in a principled and transferable manner, rather than merely adapting existing adversarial methods.
Appendix B Preliminary
B.1 Image Generation via DDPM
Denoising Diffusion Probabilistic Models (DDPM) [42] are generative models that refine images by reversing a diffusion process. Pretrained DDPMs have been applied to unconditional image generation of faces [37], CIFAR-10 [52, 53], and ImageNet [41, 54] from Gaussian distributions. In the forward diffusion process, Gaussian noise is progressively added to a data sample $x_{0}$ over $T$ time steps, generating a sequence $\{x_{t}\}_{t=1}^{T}$.
$q(x_{t}\mid x_{t-1})=\mathcal{N}(x_{t};\sqrt{\alpha_{t}}x_{t-1},(1-\alpha_{t})I),$ (23)
where $\alpha_{t}=1-\beta_{t}$ and $\beta_{t}$ is a predefined noise variance schedule. The marginal distribution of $x_{t}$ is given by:
$q(x_{t}\mid x_{0})=\mathcal{N}(x_{t};\sqrt{\bar{\alpha}_{t}}x_{0},(1-\bar{\alpha}_{t})I),$ (24)
where $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$. The reverse process aims to denoise $x_{T}\sim\mathcal{N}(0,I)$ back to a realistic image $x_{0}$ using a learned model $p_{\theta}$:
$p_{\theta}(x_{t-1}\mid x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}x_{t},t)).$ (25)
For image generation, the pretrained DDPM learns a denoising function $\epsilon_{\theta}$, parameterized by a neural network:
$\epsilon_{\theta}(x_{t},t)\approx\epsilon\sim\mathcal{N}(0,I),$ (26)
optimized through the objective:
$\mathcal{L}_{\text{DDPM}}=\mathbb{E}_{x_{0},\epsilon,t}\left[\|\epsilon-\epsilon_{\theta}(x_{t},t)\|^{2}\right].$ (27)
This noise-aware iterative refinement process enables high-fidelity image generation.
We define the final denoised image $x_{0}$ as the reconstructed image $\hat{x}$, applying a pretrained DDPM as a generator $G(\cdot)$ to an initial Gaussian noise sample $x_{G}$:
$\hat{x}=G(x_{G}).$ (28)
Here, $x_{G}$ is a latent code sampled from a Gaussian distribution, which $G(\cdot)$ refines into a high-fidelity image $\hat{x}$.
B.2 Open-Set Face Recognition
Open-set face recognition systems [4, 5, 6, 7] map facial images to embeddings and classify image pairs as the same identity if their similarity exceeds a threshold, typically chosen to balance false positives or minimize the equal error rate.
Let $F(\cdot)$ be the embedding function of a face recognition model that maps a facial image $x$ to a $d$-dimensional feature embedding $z$, encoding identity-specific attributes while ensuring robustness to variations in lighting, pose, and expression. Formally:
$z=F(x),\quad z\in\mathbb{R}^{d},$ (29)
where $d$ is determined by the model’s architecture and training process.
Feature embeddings facilitate identity verification or recognition by measuring similarity using metrics such as cosine similarity or Euclidean distance. Given two embeddings, $z_{1}=F(x_{1})$ and $z_{2}=F(x_{2})$, their similarity score is computed as:
$S(z_{1},z_{2})=\frac{z_{1}\cdot z_{2}}{\|z_{1}\|\|z_{2}\|},$ (30)
for cosine similarity. If $S(z_{1},z_{2})\geq\tau_{F}$, the images are classified as belonging to the same identity. In this work, $\tau_{F}$ denotes the predefined similarity threshold set at the minimum equal error rate.
TABLE XIV: Key notations and their corresponding definitions.
     Notation      Definition      Remark      Reference
     $F(\cdot)$      Embedding function (Face recognition)      Sec. 2.3
     $x^{T}$      Target facial image      Unknown to Attackers
     $z^{T}$      Target embedding, transformed from the target face      Known to Attackers
     $\hat{x}$      Reconstructed image      Attack Output
     $\hat{z}$      Feature embedding of the reconstructed image
     $G(\cdot)$      Generative function (DDPM)      Pretrained
     $x_{G}$      Latent code, drawn from a random Gaussian distribution      Attack Input
     $x^{\prime}_{G}$      Manipulated latent code
     $\delta$      Adversarial perturbations on the latent code
     $||\cdot||_{p}$      $L_{p}$-norm
     $\epsilon$      Perturbation magnitude      Attack Setting
     $S(\cdot,\cdot)$      Similarity function (cosine)
     $\mathcal{L}$      Objective function
     $\tau_{F}$      Similarity decision threshold
     $K(\cdot)$      Gaussian normality test function ($K^{2}$ test)      Sec. 2.4.1
     $p_{K}$      Gaussian normality (the probability of following a normal distribution)
     $\tau_{K}$      Threshold of Gaussian normality      Attack Setting
     $D(\cdot)$      Face detection function (MTCNN)      Sec. 2.4.2
     $p_{D}$      Face detection confidence score
     $\tau_{D}$      Threshold of detection confidence      Attack Setting
     $V$      Volume of reliable latent codes (Step (a))      Attack Setting      Sec. 2.5
     $N$      Top $N$ selection (Step (b))      Attack Setting
     $Q$      Query Efficiency
     $t_{max}$      Maximum iterations per adversarial attack      Attack Setting      Sec. 2.6
     $Q_{max}$      Maximum number of queries (only for black-box attacks)      Attack Setting
     $\tau_{A}$      Attack threshold (sufficient similarity)      Attack Setting
[…]
     $\mathbb{I}(\cdot)$      Indicator function      Sec. 4
     $I$      Total number of attack samples
     $x^{j\neq T}$      Facial images distinct from the target, associated with the same identity
     $J$      Total number of $x^{j\neq T}$ for each identity
Figure 15: The algorithm of the proposed two-stage latent code generation process. Initially, D’Agostino’s $K^{2}$ test $K(\cdot)$ [43, 44, 45] evaluates a randomly generated latent code $x_{G}$, retaining those that exhibit Gaussian normality with a $p_{K}$ value exceeding the threshold $\tau_{K}$. Subsequently, MTCNN $D(\cdot)$ [46] assesses whether the latent code can generate a facial image $\hat{x}$, with face detection confidence $p_{D}$ surpassing the threshold $\tau_{D}$.
Figure 16: Algorithm for selecting the top $N$ latent codes based on the highest embedding similarity between their reconstructions and the target.
Figure 17: Algorithm for the proposed Ranked Adversary in latent code manipulation. This process sequentially optimizes the top $N$ initial latent codes, ranked in Step (b), through adversarial manipulation to achieve the objective defined in Eq. 7 using the objective function in Eq. 6. The process concludes once $\mathcal{L}_{n}\geq\tau_{A},\ n=1,\dots,N$ for any manipulated latent code. If no code meets this criterion, the reconstruction with the highest $\mathcal{L}_{n}$ is selected as the final output.
B.3 D’Agostino’s K-Square Test
Given a random latent code $x_{G}$, the $K^{2}$ test function $K(\cdot)$ [43, 44, 45] quantifies deviations from normality based on skewness and kurtosis, producing a probability value:
$p_{K}=K(x_{G})=1-\Psi_{2}(Y_{1}^{2}+Y_{2}^{2}),$ (31)
where $Y_{1}$ and $Y_{2}$ are the standardized skewness and kurtosis statistics:
$Y_{1}=\frac{b_{1}-\mu_{1}}{\sigma_{1}},\quad Y_{2}=\frac{b_{2}-\mu_{2}}{\sigma_{2}}.$ (32)
Here, $b_{1}$ and $b_{2}$ represent the sample skewness and excess kurtosis, while $\mu_{1},\sigma_{1},\mu_{2},$ and $\sigma_{2}$ denote their respective means and standard deviations under normality. The function $\Psi_{2}(\cdot)$ is the cumulative distribution function of the chi-squared distribution with two degrees of freedom.
B.4 Face Detection via MTCNN
Multi-Task Cascaded Convolutional Neural Networks (MTCNN) [46] uses a three-stage cascaded architecture: (i) The Proposal Network (P-Net) scans the image across multiple scales to generate candidate face regions. (ii) The Refinement Network (R-Net) filters false positives and refines bounding boxes. (iii) The Output Network (O-Net) further refines detections and predicts five facial landmarks. The final detection confidence is obtained through:
$p_{D}=D(\hat{x})=f_{\text{O-Net}}(f_{\text{R-Net}}(f_{\text{P-Net}}(\hat{x}))).$ (33)
Appendix C Algorithms of DiffUMI
The step-by-step algorithm of the proposed model inversion attack, DiffUMI, is detailed in Figs. 15, 16 and 17. The key notations and their corresponding definitions are summarized in Tab. XIV.
Appendix D More Results in Terms of Type I Accuracy
In Sec. 4, we introduce Type I and Type II accuracy as metrics for assessing the effectiveness of privacy attacks. This study primarily focuses on Type II accuracy, which quantifies the similarity between reconstructed facial images and images from the target identity, excluding the target image itself. Type II accuracy offers a more rigorous evaluation by mitigating the risk of overfitting to the target image, whose embedding is used as a reference during the attack. For further validation, we also report Type I accuracy results, which reflect the strength of the inversion attack and highlight potential privacy vulnerabilities. Specifically, Type I accuracy is the rate at which the reconstructed image $\hat{x}$ matches the target face $x^{T}$ in the feature space of the face recognition model $F(\cdot)$:
$\text{Type I}=\quad\frac{\sum_{i=1}^{I}\mathbb{I}\left(S(F(\hat{x_{i}}),F(x^{T}_{i}))\geq\tau_{F}\right)}{I},$ (34)
where $\mathbb{I}(\cdot)$ and $S(\cdot,\cdot)$ represent the indicator and cosine similarity functions, respectively. $I$ is the total number of attack samples, and $\tau_{F}$ is the similarity threshold for face recognition.
The results in Tabs. XV and XVI reinforce our main paper’s conclusion that all tested face recognition models are vulnerable to our privacy attack. Among them, FaceNet, the oldest standard model, exhibits the highest resistance, underscoring the limitations of existing privacy-preserving techniques. Our approach consistently outperforms the benchmark MAP2V attack [9], achieving higher Type I accuracy across all scenarios. Notably, Type I accuracy exceeds Type II accuracy by 3.83% to 15.79%, demonstrating that our primary evaluation metric is more rigorous and comprehensive.
TABLE XV: Type I accuracy (%) of DiffUMI across four face recognition models.
Dataset Target Test Model Avg.
Model FaceNet ArcFace DCTDP PartialFace
CelebA FaceNet 100.0 96.7 97.9 94.9 97.38
ArcFace 99.9 100.0 100.0 99.9 99.95
DCTDP 99.3 99.5 100.0 99.9 99.68
PartialFace 97.7 99.2 99.6 100.0 99.13
LFW FaceNet 100.0 87.0 93.4 80.8 90.30
ArcFace 96.0 100.0 99.6 98.6 98.55
DCTDP 96.7 99.7 100.0 97.8 98.55
PartialFace 90.9 98.7 96.2 100.0 96.45
  • Gray cells indicate cases where the target and test models are identical. Green and Red highlight the most and least secure models, respectively, based on the lowest and highest Type I accuracy.
TABLE XVI: Type I accuracy (%) of our DiffUMI across four face recognition models, compared to the benchmark attack MAP2V [9].
Dataset Attack Target Model
FaceNet ArcFace DCTDP PartialFace
CelebA Original1 100.00 100.00 100.00 100.00
Random2  4.40  1.30  4.20 10.10
MAP2V 89.75 95.03 96.75 97.98
Ours 97.38 99.95 99.68 99.13
LFW Original1 100.00 100.00 100.00 100.00
Random2  0.70  0.00  0.20  0.50
MAP2V 90.05 88.98 92.50 90.54
Ours 90.30 98.55 98.55 96.45
  • 1 Upper bound corresponding to true target faces.
  • Bold indicates the highest performance among attack methods, with our approach exceeding MAP2V attack by up to 9.57% in Type I accuracy.